# Solvation Free Energies from Neural Thermodynamic Integration

### v2

@@ I Introduction
In this work, we apply the same approach to a scenario where both endpoint Hamiltonians are not trivial, but represent coupled and decoupled states of a solute–solvent system as illustrated in Fig. 1. We generate samples $x_{t}\sim\rho_{t}$ as the geodesic interpolation between samples $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1}$ and optimize a neural network potential to match the corresponding equilibrium potential at every $10<t<$ 1. Notably, in our experiments on chemical systems, the Lennard-Jones and electrostatic interactions are coupled simultaneously as parametrized by the neural network. The paper is structured as follows: Section II summarizes the relevant concepts that our work builds on. In Section III we argue that ideas from the machine-learning community generalizing diffusion models provide a natural framework for this setting [25, 26, 27] and describe our approach. Finally, in Section IV we validate our methodology by computing solvation free energies on several systems: ($i$) an LJ solute in an LJ solvent, and the insertion of ($ii$) a water solute and ($iii$) a methane solute in a box of water, both at atomistic resolution. All free energies show excellent agreement with reference free energies.
@@ II Background
We begin with fixing some notation. Depending on which arguments are important in the context, we will use $U_{t}(x)$ or $U(t,x)$ or simply $U$ to denote time-dependent energy/potential functions and $\rho_{t}$ for the corresponding Boltzmann probability density $\frac{1}{Z_{t}}e^{-\beta U_{t}}$, with inverse temperature $\beta=1/k_{\mathrm{B}}T$. The symbol $\nabla U$ is used for the spatial gradient. The isotropic Gaussian distribution, centered at the origin, with variance $\sigma^{2}$, is written as $\mathcal{N}_{\sigma^{2}}$ and the convolution of two functions, $f$ and $g$ as $f\star g$.
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under $\rho_{\lambda}=Z_{\lambda}^{-1}e^{-\beta U_{\lambda}}$. To numerically estimate the value of the last integral one needs an interpolating potential function $U_{\lambda}$ and samples from the corresponding interpolating equilibrium distributions $\rho_{\lambda}$. Usually one first fixes an interpolation $U_{\lambda}$ and performs Monte Carlo sampling or molecular dynamics (MD) simulations at intermediate $\lambda$ values to obtain the estimate
$\Delta\hat{F}_{0\rightarrow 1}=\frac{1}{N+1}\sum_{\lambda_{i}}\,\mathbb{E}_{x\sim\rho_{\lambda_{i}}}\big{[}\partial_{\lambda}U_{\lambda}(x)\big{]},$ (3)
where the sum is over $\lambda_{i}\in\{0,1/N,2/N,...,1\}$.
@@
Máté et al. [24] proposed to parametrize a time-dependent energy function $U_{t}$ and use score matching (SM) to align the force $\nabla U_{t}$ with the scores of increasingly noised versions of the target Boltzmann distribution $e^{-\beta U_{0}}/Z_{0}$. As the noise level $\sigma(t)$ starts small at $t=0$, and is large at $t=1$, the trained $U_{t}$ provides an interpolating potential between the target and a tractable prior. Since samples from the corresponding Boltzmann distributions $e^{-\beta U_{t}}\approx\mathcal{N}_{\sigma_{t}^{2}}\star e^{-\beta U_{0}}$ are also (approximately) available, TI can be performed using $U_{t}$ to estimate the free-energy difference between a trivial and a non-trivial distribution. In this case, these were matched to the ideal gas and the Boltzmann distribution of the fully coupled LJ liquids, so as to associate the resulting free energy to the coupling of all the interactions. Traditional approaches to solvation free-energy calculations often consist of first fixing the interpolation in the space of potentials (e.g., linear coupling), to subsequently generate samples from Monte Carlo sampling or molecular dynamics simulations [15]. Here instead, we first fix the sampling procedure and learn the interpolating family of potentials.
@@
A possible issue arises for a domain that is disconnected or a target distribution that is multimodal: the additive constant becomes local. Thus, strictly speaking, our model does not learn a function $U_{t}$ such that $e^{-\beta U_{t}}$ is globally proportional to $\rho_{t}$, but only locally proportional. In other words, it does not preserve the relative weights of the modes, but does preserve the relative likelihoods within the modes. Song and Ermon [29] solve this by learning a family of models with increasing noise levels, making use of the fact that for high perturbations the target distribution becomes connected. Alternatively, Gutmann and Hyvärinen [30] overcome the issue by contrasting the data with a second distribution of known density. In the present work, the enforced boundary conditions alleviate the issue: All offsets introduced to the relative weighting of the modes also integrate to zero over the time interval $[0,1]$. See Figure 2 for a one-dimensional illustration.
@@ III Method
The goal of this work is to extend the neural TI framework to the estimation of free-energy differences between two non-trivial Hamiltonians. To this end, we consider a pair of potential functions, $U_{0},U_{1}:\mathcal{M}\rightarrow\mathbb{R}$ on a manifold $\mathcal{M}$ representing the configuration space of some system. In our experiments, we model particle systems and rigid molecules in a periodic box. For the particle system $\mathcal{M}$ will be the $3N$-torus, $\mathbb{T}^{3N}$ representing the positions of $N$ particles in a 3-torus. For the rigid molecular system, the configuration of each molecule is given by its position in $\mathbb{T}^{3}$ and its orientation, thus the configuration space of the full system is $\mathcal{M}=[\mathbb{T}^{3}\times SO(3)]^{N}$.
@@
De Bortoli et al. [35] recently proposed target score matching (TSM), an alternative SM objective that is well-suited for our application. TSM relies on having access to target score functions at the endpoints, $\nabla\log\rho_{i}=-\nabla U_{i}$ for $i\in\{0,1\}$. Decomposing $\nabla\log\rho_{t}$, as
@@
In our experiments we combine these two objectives and work with the following target score matching objective,
@@
where $U_{\mathrm{solvent}},U_{\mathrm{solute}}$ and $U_{\mathrm{solute-solvent}}$ represent the interactions within the solvent, within the solute and cross interactions between the two components, respectively. In Section IV we consider two experimental setups, ($i$) an LJ solute included in an LJ solvent and the inclusion of a rigid solute molecule in a solvent of rigid water molecules. We impose periodic boundary conditions, leading to configuration spaces $\mathbb{T}^{3}$ and $[\mathbb{T}^{3}\times SO(3)]^{N}$, respectively. The solute is rigid in both cases, implying $U_{\mathrm{solute}}\equiv 0$. As described below, most properties of these systems are shared. When discussing the differences between them we refer to the two setups by LJ-experiment and hydration-experiment.
@@
where $r$ is the distance between the particles and $(\epsilon,\sigma)$ denote the energy and length scales of the LJ interactions. When $a=0$, no softening is applied, and larger values of $a$ mean larger softening deformation of the original Lennard-Jones interactions. Note that equation (15) corresponds to evaluating the unsoftened Lennard-Jones potential at $\sqrt{r^{2}+a\sigma^{2}}$. Interactions between different species, $A$ and $B$, of particles are computed by the Lorentz-Berthelot rules [37, 38] $(\varepsilon_{AB}=\sqrt{\varepsilon_{A}\varepsilon_{B}},\sigma_{AB}=(\sigma_{A}+\sigma_{B})/2)$.
@@
where $U_{t}^{\theta}(x)$ is a neural network with trainable parameters. We enforce the boundary conditions of the softening $t\rightarrow a_{t}^{A},a^{AB}_{t},a_{t}^{B}$ and coupling functions $t\rightarrow b_{t}^{A},b^{AB}_{t},b_{t}^{B}$ by the parametrization
@@
where the $t\mapsto\alpha_{t}^{*},t\mapsto\beta_{t}^{*}$ functions are learned by a small multilayer perceptron. These functional forms ensure that at $t=0$ and $t=1$ the solute and solvent interactions are fully coupled and no softening is applied to them, while the solvent–solute interactions are fully decoupled at $t=0$ and coupled at $t=1$. We predict such a set of parameters for each interaction separately.
@@
$\mathcal{L}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x_{t}\sim\rho_{t}}\Big{[}\mathcal{L}_{\mathrm{SM}}+\kappa(\partial_{t}U_{t}(x_{t}))^{2}\Big{]},$ (26)
where $U_{t}$ depends on $\theta$ through equation 20 and $\mathcal{L}_{\mathrm{SM}}\in\{\mathcal{L}_{\mathrm{DSM}},\mathcal{L}_{\mathrm{TSM}}\}$. Since $\partial_{t}U_{t}$ is the very thing we are interested in measuring, we use a small relative weight for its regularization in our experiments, $\kappa=10^{-10}$. The goal is not to minimize the $L_{2}$-norm of $\partial_{t}U_{t}$, but rather to prevent it from diverging.
@@ IV Results
We compare our estimates with TI using evenly spaced time-slices along the time-interval $[0,1]$. We perform 51 simulations in total at $t\in\{0.00,0.02,0.04,...,1.00\}$, and find that the solvation free energy lies between $-21k_{B}T$ and $-19k_{B}T$. The rolling estimate of neural TI converges to approximately $-19.5k_{B}T$. The time evolution of these values are shown in the left subplot of Figure 3. The solvation free energy estimate uses data only from the two endpoints and shows significantly lower variance than the associated standard TI computation, that requires between 5 and 50 interpolating timesteps. The right subplot of Figure 3 displays the learned softening schedule of the solvent–solvent and solute–solvent Lennard-Jones interactions. The learned softening functions show that our interpolation does not correspond to the usual linear interpolation $(1-t)U_{0}+tU_{1}$, even the interactions within the solvent are softened and decoupled along the trajectory.
Figure 3: Solute solvation of a Lennard-Jones fluid. Left: Solvation free-energy estimates by TI and neural TI as a function of number of reference simulations and training steps, respectively. Center: Detailed view at later training stages. Softening parameters of the LJ-interactions after training (right).
@@
Given that the two solutes are of comparable size and complexity, we may assume that two systems, and in particular their solute–solvent coupling process, share some similarities. In the ideal case, a model trained on one of the solutes could give a reasonable prediction when evaluated on the other. We test this with the trained models and confirm that this is indeed the case: both models learn features that are useful when modeling the coupling of the other solute and predict the hydration free energy of the unseen solute within a $4k_{B}T$ accuracy. We attribute this transferability to our ansatz (equation 20), which enforces the target Hamiltonians, along with the contribution of solvent-solvent interactions. The learned softening of the endpoints and the accurate modeling of solvent-solvent interactions throughout the coupling process are likely key factors underlying the observed transferability.
All numerical values are reported in Table 1.
@@ V Conclusion
The method is verified by accurately estimating the solvation free energy of a Lennard-Jones solute in a Lennard-Jones solvent using only data from the endpoint simulations, matching the accuracy of standard TI, which typically requires tens of intermediate simulations. We also predict the hydration free energies of rigid water and methane, showing good agreement with experimental values, using a simple three-body potential. Future work could explore introducing bonded interactions into the system, which would eliminate the need for a non-Euclidean configuration space but would require handling forces acting on different length scales. Finally, we observed promising transferability of the learned potential between these two solutes.

### v3

@@ I Introduction
In this work, we apply the same approach to a scenario where both endpoint Hamiltonians are not trivial, but represent coupled and decoupled states of a solute–solvent system as illustrated in Fig. 1. This in particular means that we can no longer rely on the diffusion process for interpolation, but need to choose a different way of generating intermediate samples. To do this, we take inspiration from recent generalization of diffusion models [25, 26] and construct samples $x_{t}\sim\rho_{t}$ as the geodesic interpolation between samples $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1}$. Given these intermediate samples, the approach of neuralTI carries over to the current setting and we can optimize a neural network potential to match the equilibrium potential of $\rho_{t}$ at every $10<t<$ 1. In our experiments on chemical systems, we model both the Lennard-Jones and electrostatic interactions and couple them simultaneously as parametrized by the neural network.
The paper is structured as follows: Section II summarizes the relevant concepts that our work builds on. In Section III we argue that ideas from the machine-learning community generalizing diffusion models provide a natural framework for this setting [25, 26, 27] and describe our approach. Finally, in Section IV we validate our methodology by computing solvation free energies on several systems: ($i$) an LJ solute in an LJ solvent, and the insertion of ($ii$) a water solute and ($iii$) a methane solute in a box of water, both at atomistic resolution. All free energies show excellent agreement with reference free energies.
@@ II Background
We begin with fixing some notation. Depending on which arguments are important in the context, we will use $U_{t}(x)$ or $U(t,x)$ or simply $U$ to denote time-dependent potential functions and $\rho_{t}(x)$ for the corresponding Boltzmann probability density $\frac{1}{Z_{t}}e^{-\beta U_{t}(x)}$, with inverse temperature $\beta=1/k_{\mathrm{B}}T$. The symbol $\nabla U$ is used for the spatial gradient. The isotropic Gaussian distribution, centered at the origin, with variance $\sigma^{2}$, is written as $\mathcal{N}_{\sigma^{2}}$ and the convolution of two functions, $f$ and $g$ as $f\star g$.
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expectation value of $\partial_{\lambda}U_{\lambda}$ under $\rho_{\lambda}=Z_{\lambda}^{-1}e^{-\beta U_{\lambda}}$. To numerically estimate the value of the last integral one needs an interpolating potential function $U_{\lambda}$ and samples from the corresponding interpolating equilibrium distributions $\rho_{\lambda}$. Usually one first fixes an interpolation $U_{\lambda}$ and performs Monte Carlo sampling or molecular dynamics (MD) simulations at intermediate $\lambda$ values to obtain the estimate
$\Delta\hat{F}_{0\rightarrow 1}=\frac{1}{N}\sum_{\lambda_{i}}\,\mathbb{E}_{x\sim\rho_{\lambda_{i}}}\big{[}\partial_{\lambda}U_{\lambda}(x)\big{]},$ (3)
where the sum is over $\lambda_{i}\in\{0,1/N,2/N,...,(N-1)/N\}$.
@@
Máté et al. [24] proposed to parametrize a time-dependent energy function $U_{t}$ and use score matching (SM) to align the force $\nabla U_{t}$ with the scores of increasingly noised versions of the target Boltzmann distribution $e^{-\beta U_{0}}/Z_{0}$. As the noise level $\sigma(t)$ starts small at $t=0$, and is large at $t=1$, the trained $U_{t}$ provides an interpolating potential between the target and a tractable prior. Since samples from the corresponding Boltzmann distributions $e^{-\beta U_{t}}=\mathcal{N}_{\sigma_{t}^{2}}\star e^{-\beta U_{0}}$ are also (approximately) available, TI can be performed using $U_{t}$ to estimate the free-energy difference between a trivial and a non-trivial distribution. The prior and data densities were matched to the ideal gas and the Boltzmann distribution of the fully coupled LJ liquids, so as to associate the resulting free energy to the coupling of all the interactions. Traditional approaches to solvation free-energy calculations often consist of first fixing the interpolation in the space of potentials (e.g., linear coupling), to subsequently generate samples from Monte Carlo sampling or molecular dynamics simulations [15]. Here instead, we first fix the sampling procedure and learn the interpolating family of potentials.
@@
A possible issue arises for a domain that is disconnected or a target distribution that is multimodal: the additive constant becomes local. Thus, strictly speaking, our model does not learn a function $U_{t}$ such that $e^{-\beta U_{t}}$ is globally proportional to $\rho_{t}$, but only locally proportional. In other words, it does not preserve the relative weights of the modes, but does preserve the relative likelihoods within the modes. Song and Ermon [29] solve this by learning a family of models with increasing noise levels, making use of the fact that for high perturbations the target distribution becomes connected. Alternatively, Gutmann and Hyvärinen [30] overcome the issue by contrasting the data with a second distribution of known density. In the present work, the enforced boundary conditions alleviate the issue: All offsets introduced to the relative weighting of the modes also integrate to zero over the time interval $[0,1]$. See Figure 2 for a one-dimensional illustration. While mathematically equivalent, from a numerical perspective, interpolations that yield lower variance of $\partial_{t}U_{t}$ are preferred for the TI-estimation of the free-energy difference. Whether the training converges to such a potential likely depends on the initialization scheme, the training dynamics and the explicit regularization of the variance of $\partial_{t}U_{t}$.
@@ III Method
The goal of this work is to extend the neural TI framework to the estimation of free-energy differences between two non-trivial Hamiltonians. To this end, we consider a pair of potential functions, $U_{0},U_{1}:\mathcal{M}\rightarrow\mathbb{R}$ on a manifold $\mathcal{M}$ representing the configuration space of some system.
In this general setting, without further assumptions on the system, we can outline our proposed method.
  1. 1.
    Choose a parametric family of interpolations between $U_{0}$ and $U_{1}$. We can always use $U(t,x)=(1-t)U_{0}(x)+tU_{1}(x)+t(1-t)U^{\theta}(t,x)$, where $U^{\theta}(t,x)$ is a parametric function, usually a neural network, with parameters $\theta$. The only requirement here is that the boundary conditions at $t\in\{0,1\}$ are satisfied.
  2. 2.
    Define the sampling process of $\rho_{t}$—in other words, a way to construct samples $x_{t}\sim\rho_{t}$ given samples $(x_{0},x_{1})\sim(\rho_{0},\rho_{1})$. This highly depends on the geometry of $\mathcal{M}$. As long as $\mathcal{M}$ can be equipped with the structure of a Riemannian manifold, i.e., a metric $g\in C^{\infty}(\mathcal{M},\text{Symm}^{2}(T^{*}\mathcal{M}))$, the geodesic interpolation with respect to $g$ is a straightforward choice.
  3. 3.
    Given the two above-mentioned constructions, what remains is to optimize $U^{\theta}(t,x)$ until it approximates the equilibrium potential of $\rho_{t}$. In this work we do this by score matching.
In our experiments, we focus on solvation free energies by modeling the intermolecular interactions between solute and solvent using common molecular-mechanics terms, except that we freeze intramolecular contributions by working with rigid molecules. All systems are embedded in a periodic box. For the particle system $\mathcal{M}$ will be the $3N$-torus, $\mathbb{T}^{3N}$ representing the positions of $N$ particles in a 3-torus. For the rigid molecular system, the configuration of each molecule is given by its position in $\mathbb{T}^{3}$ and its orientation, thus the configuration space of the full system is $\mathcal{M}=[\mathbb{T}^{3}\times SO(3)]^{N}$.
@@
De Bortoli et al. [35] recently proposed target score matching (TSM), an alternative SM objective that is well-suited for our application. TSM relies on having access to target score functions at the endpoints, $\nabla\log\rho_{0}=-\nabla U_{0}$ and $\nabla\log\rho_{1}=-\nabla U_{1}$. Decomposing $\nabla\log\rho_{t}$, as
@@
In our experiments we combine these two terms and work with a target score matching objective that always uses the closer endpoint of the $[0,1]$ interval,
@@
where $U_{\mathrm{solvent}},U_{\mathrm{solute}}$ and $U_{\mathrm{solute-solvent}}$ represent the interactions within the solvent, within the solute and the cross interactions between the two components, respectively. In Section IV we consider two experimental setups, ($i$) an LJ solute included in an LJ solvent and ($ii$) the inclusion of a rigid solute molecule in a solvent of rigid water molecules. We impose periodic boundary conditions, leading to configuration spaces $\mathbb{T}^{3}$ and $[\mathbb{T}^{3}\times SO(3)]^{N}$, respectively. The solute is rigid in both cases, implying $U_{\mathrm{solute}}\equiv 0$. As described below, most properties of these systems are shared. When discussing the differences between them we refer to the two setups by LJ experiment and hydration experiment.
@@
where $r$ is the distance between the particles and $(\epsilon,\sigma)$ denote the energy and length scales of the LJ interactions. When $a=0$, no softening is applied, and larger values of $a$ mean larger softening deformation of the original Lennard-Jones interactions. Note that equation (15) corresponds to evaluating the unsoftened Lennard-Jones potential at $\sqrt{r^{2}+a\sigma^{2}}$. The parameters of the LJ interactions between different species, $A$ and $B$, of particles are computed by the Lorentz-Berthelot rules [37, 38], $\varepsilon_{AB}=\sqrt{\varepsilon_{A}\varepsilon_{B}},\sigma_{AB}=\frac{1}{2}(\sigma_{A}+\sigma_{B})$.
@@
where $U_{t}^{\theta}(x)$ is a neural network with trainable parameters. We enforce the boundary conditions of the softening $t\rightarrow a_{t}^{A},a^{AB}_{t},a_{t}^{B}$ and coupling functions $t\rightarrow b_{t}^{A},b^{AB}_{t},b_{t}^{B}$ with the parametrization
@@
where the $t\mapsto\alpha_{t}^{*},t\mapsto\beta_{t}^{*}$ functions are learned by a multilayer perceptron with two hidden layers, each containing 256 hidden units, and swish activations [41]. These functional forms ensure that at $t=0$ and $t=1$ the solute and solvent interactions are fully coupled and no softening is applied to them, while the solvent–solute interactions are fully decoupled at $t=0$ and coupled at $t=1$. We predict such a set of parameters for each interaction separately.
@@
$\mathcal{L}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x_{t}\sim\rho_{t}}\Big{[}\lambda_{t}\mathcal{L}_{\mathrm{SM}}+\kappa(\partial_{t}U_{t}(x_{t}))^{2}\Big{]},$ (26)
where $U_{t}$ depends on $\theta$ through equation 20, $\mathcal{L}_{\mathrm{SM}}\in\{\mathcal{L}_{\mathrm{DSM}},\mathcal{L}_{\mathrm{TSM}}\}$, and $\lambda_{t}$ is the time-dependent weighting of the score matching term. Since $\partial_{t}U_{t}$ is the very thing we are interested in measuring, we use a small relative weight for its regularization in our experiments, $\kappa\in\{10^{-8},10^{-10}\}$. The goal is not to minimize the $L_{2}$-norm of $\partial_{t}U_{t}$, but rather to prevent it from diverging.
@@ IV Results
We compare our estimates with TI using evenly spaced time-slices along the time-interval $[0,1]$. We perform 51 simulations in total at $t\in\{0.00,0.02,0.04,...,1.00\}$, and find that the solvation free energy lies between $-21k_{B}T$ and $-19k_{B}T$. The rolling estimate of neural TI converges to approximately $-19.5k_{B}T$. The time evolution of these values are shown in the left subplot of Figure 3. The solvation free energy estimate uses data only from the two endpoints and shows significantly lower variance than the associated standard TI computation, that requires between 5 and 50 interpolating timesteps. The right subplot of Figure 3 displays the learned softening schedule of the solvent–solvent and solute–solvent Lennard-Jones interactions. The learned softening functions show that our interpolation does not correspond to the usual linear interpolation $(1-t)U_{0}+tU_{1}$, even the interactions within the solvent are softened and decoupled along the trajectory. Consistent with our earlier work on neural TI [24], we find that neural TI and traditional TI converge in a comparable amount of time.
Figure 3: Solute solvation of a Lennard-Jones fluid. Left: Solvation free-energy estimates by TI and neural TI as a function of number of reference simulations and training steps, respectively. Center: Detailed view at later training stages. Softening parameters of the LJ interactions after training (right).
@@
Given that the two solutes are of comparable size and complexity, we may assume that two systems, and in particular their solute–solvent coupling process, share some similarities. In the ideal case, a model trained on one of the solutes could give a reasonable prediction when evaluated on the other. We test this with the trained models and confirm that this is indeed the case: both models learn features that are useful when modeling the coupling of the other solute and predict the hydration free energy of the unseen solute within a $4k_{B}T$ accuracy. We attribute this transferability to our ansatz (equation 20), which enforces the target Hamiltonians, along with the contribution of solvent-solvent interactions. The learned softening of the endpoints and the accurate modeling of solvent-solvent interactions throughout the coupling process are likely key factors underlying the observed transferability. All numerical values are reported in Table 1.
@@ V Conclusion
The method is verified by accurately estimating the solvation free energy of a Lennard-Jones solute in a Lennard-Jones solvent using only data from the endpoint simulations, matching the accuracy of standard TI, which typically requires tens of intermediate simulations. We also predict the hydration free energies of rigid water and methane, showing good agreement with experimental values, using a simple three-body potential. Importantly, we do not observe or claim a computational speedup compared to traditional TI. We do however see promise for a computational advantage by learning an interpolating potential that is transferable between different systems. We observe a simple instance of such a transferable potential between water and methane in our hydration experiment (Table 1).
Future work could scale the method to more complex systems. The main limitation of the current approach is the rigidity assumption of the molecules. Overcoming this requires introducing bonded interactions, which removes the need for the $SO(3)$ factors in the configuration space but also introduces forces acting on different length scales, potentially posing numerical challenges. Additionally, it would be of interest to control the variance of the TI estimate (eq. 12) without explicitly regularizing it (eq. 26).
Code availability
Our implementation is available at the following repository: https://github.com/balintmate/neuralTI-solvation.